a.Suppose a market research company finds that at a price of p = $12, they would sell x = 54 tiles each month. If they lower the price to p = $6, then more people would purchase the tile, and they can expect to sell x = 60 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p). Show how you arrived at your answer.
Answer by stanbon(75887)

a.Suppose a market research company finds that at a price of p = $12, they would sell x = 54 tiles each month. If they lower the price to p = $6, then more people would purchase the tile, and they can expect to sell x = 60 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p).
----
Two points: (12,54) and (6,60)
---
slope = (60-54)/(6-12) = 6/-6 = -1
----------
Form p = mx + b
60 = -1(6) + b
b = 66
-----
Equation:
p(x) = -x + 66
====================
Cheers,
Stan H.
=================